OpenStudy (anonymous):

Determine dy/dx: y=xe^-x^2. I got y'=-xe^-x^2+e^-x^2. Am i correct?

OpenStudy (anonymous):

\[\frac{d}{dx}xe^{-x^2}\]?

OpenStudy (anonymous):

then no the answer is wrong. should be \[e^{-x^2}-2x^2e^{-x^2}\]

OpenStudy (anonymous):

via product rule and chain rule

OpenStudy (anonymous):

\[\frac{d}{dx}e^{-x^2}=-2xe^{-x^2}\]

OpenStudy (anonymous):

f(x)=x*e^u\[ dy/dx=e^u + x *u \prime*e^u\]